# DiSHA: Dimension-Sharding Adaptation with Fast Convergence and Fast Computation

### v6

@@ Abstract
Low-Rank Adaptation (LoRA) leverages the low intrinsic rank of weight updates in Large Language Models (LLMs), establishing a Parameter-Efficient Fine-Tuning (PEFT) paradigm. However, LoRA suffers from slow convergence. We introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to unlock lower intrinsic ranks and faster convergence by default. Within DiSHA’s design space, we propose Block Affine Adaptation (Bone), a computationally efficient structure that delivers both high performance and efficiency. While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation Adaptation (BAT), a nonlinear variant of DiSHA. BAT introduces nonlinearity by combining trainable matrices with original weight shards in a nonlinear manner, inducing nonlinearity in matrix updates without introducing additional parameters. Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both NLG and NLU tasks, with significantly improved computational efficiency. Further analysis demonstrates that BAT enhances model capabilities by leveraging its nonlinear design.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is computationally costly. As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)Xu et al. (2023) techniques and quantization methods have emerged to reduce the training costs of large models. Low-Rank Adaptation(LoRA) Hu et al. (2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning. The figure  3 illustrates the structure visualization.
@@
However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). These methods have all brought unexpected improvements. By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
@@
In this work, we first propose a novel framework called Dimension-Sharding Adaptation (DiSHA) to design PEFT with lower ranks and faster convergence. In DiSHA, the pre-trained weights $W_{0}$ are frozen and divided into multiple shards during training, with all shard weights sharing a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters. The design space of DiSHA is vast, allowing pre-trained weights to be divided into dimensions of arbitrary sizes 2. Computational efficiencies vary among design choices. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments. Of course, we also observed that certain subsets of the DiSHA design space exhibit limitations: all shard weight updates are colinear, which could potentially hinder model performance. To address this potential issue, we propose the Bat method. Without introducing additional trainable matrices, Bat uses the trainable matrix to interact with the original weight shards, transforming the linear updates between the shards into non-linear updates. This approach enhances the model’s capability while maintaining efficiency.
Figure 2: partition of weight matrix $W_{0}$.
Our extensive evaluation shows that Bone demonstrates excellent performance on both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks, significantly surpassing LoRA and its variants. Moreover, as shown in Figure  5, Bone also achieves a significantly faster convergence rate compared to PiSSA. This demonstrates the feasibility of the DiSHA framework. Furthermore, Bone’s unique design not only resolves the unification challenges of the DiSHA framework, allowing seamless adaptation across different LLM architectures such as LLaMA, RWKV, and Mistral but also enables it to significantly outperform LoRA in both computational efficiency and memory usage. In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). Based on the comparative experiments between Bone and Bat, it is evident that the non-linear updates of weight shards enhance the model’s performance ceiling.
@@
2. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
3. We identified a potential issue in the DiSHA framework: the updates of weight shards are colinear. To address this, we improved upon Bone and proposed Bat, which transforms weight shard updates into non-linear ones, thereby enhancing the model’s performance ceiling.
@@ 2 Related Works
The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al. (2019); He et al. (2022); Wang et al. (2022); Pfeiffer et al. (2020), prefix tuningLiu et al. (2023); Li & Liang (2021), prompt tuningBrown (2020); Liu et al. (2023); Lester et al. (2021); Razdaibiedina et al. (2023); Li & Liang (2021), LoRAHu et al. (2021); Meng et al. (2024); Wang et al. (2024); Si et al. (2024), and layer-freezing methods such as LISA.
@@
LoRA, as one of the most popular PEFT techniques, has demonstrated that the updates to the weights exhibit a low intrinsic rank during adaptation. In LoRA, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters. Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $h=W_{0}x$, our modified forward pass yields:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (1)
@@
Since A and B are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process. As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial gradients, among other approaches.
@@
LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time. Therefore, we need to design novel structures to facilitate the exploration of a lower intrinsic rank.
@@ 3.1 DiSHA: Dimension-Sharding Adaptation
Hypothesis Based on the conclusions of the experimental analysis of the gradient norm  9.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower intrinsic ranks and faster convergence by default. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W=W_{0}+\Delta W=W_{0}+expand(D)$, where $D\in\mathbb{R}^{r_{1}\times r_{2}},expand(D)\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $D$ is a trainable parameter initialized to 0. Note both $W_{0}$ and $\Delta W=expand(D)$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $y=W_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
$y=W_{0}x+\Delta Wx=W_{0}x+expand(D)x$ (2)
Here, we explain the concept of $expand(D)$. As shown in the DiSHA structure in Figure  3, we divide $W_{0}$ into multiple shards along the dimension of $D$. The $D$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $W_{0}$. Finally, the updated shards are summed to update $W$.
@@ 3.2 DiSHA design space
In the previous section, we introduced the design principles of DiSHA and defined its trainable matrix $D\in\mathbb{R}^{r_{1}\times r_{2}}$. With DiSHA, you can customize the size of the trainable matrix $D$ based on the dimensions of the pre-trained weights. Since all shards of the pre-trained weights $W_{0}$ share a single $D$ for updates, as long as the expanded $expand(D)$ matches the dimensions of $W_{0}$, you can freely choose how to divide $W_{0}$ into shards. This flexibility makes the design space of DiSHA immense 2.
3.3 Bone: Block-Affine efficient update
@@
In the previous section, we introduced the DiSHA framework, a method that significantly reduces the number of trainable parameters in the model. Due to the vast design space of the DiSHA framework, it is difficult for users to accurately determine suitable values for training. A good setting of $D\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks. For example, with the rise of technologies like GQA Ainslie et al. (2023) and MQA Shazeer (2019), the matrix dimensions of self-attention modules ($W_{k},W_{v}$) vary greatly among different LLM models. At the same time, computational efficiency and resource utilization also vary depending on the design space choices within the DiSHA framework.
To address these limitations, we proposed Block Affine Adaptation (Bone) based on the DiSHA framework. Inspired by LoRA, we modified $D\in\mathbb{R}^{r_{1}\times r_{2}}$ to $D\in\mathbb{R}^{r\times d}$ or $D\in\mathbb{R}^{r\times k}$. This allows the trainable parameter count to be controlled by a single parameter $r$, making it easy to adapt to any neural network. Additionally, we observed that $expand(D)x$ not only increases memory consumption but also reduces computational efficiency. From our analysis, when $D\in\mathbb{R}^{r\times d}$, $x\in\mathbb{R}^{b\times l\times k}$, $expand(D)\in\mathbb{R}^{d\times k}$, the operation $expand(D)x$ essentially performs a summation along the $k$-dimension, where $expand(D)$ involves replicating the necessary number of shards. Clearly, we can first $sum(x)$ along the $k$dimension and then compute it directly with $D$, where $sum(x)\in\mathbb{R}^{b\times l\times r}$. This approach not only reduces memory consumption but also improves computational efficiency. The update for $W$ is shown in the figure  3, and the Bone calculation formula is as follows:
$y=W_{0}x+\Delta Wx=W_{0}x+sum(x)D$ (3)
Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
3.4 Bat: Block-Affine Transformation
We conducted extensive experiments on both NLU and NLG tasks to validate the effectiveness of Bone. It outperforms many LoRA variants and surpasses LoRA in terms of memory consumption and computational efficiency. However, Bone still has a potential issue, which stems from the shared update mechanism of the DiSHA framework. Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be colinear. This restricts weight updates and, consequently, the model’s capabilities. To address the issue of linear correlation among shard updates, our initial idea was to use a trainable coefficient matrix to control the updates of different shards. However, this approach would increase additional parameters.
To avoid introducing additional parameters, we applied a linear transformation between the original weight $W_{0}$ shards and the trainable matrix, ensuring that the updates of different shards are nonlinearly related. Based on this, we proposed the Block Affine Transformation Adaptation (Bat) method. It extends Bone by further dividing both the trainable matrix and $W_{0}$ into smaller shards, where each shard interacts with the trainable matrix $D$ through a linear transformation, achieving nonlinear independence among shard updates. The specific formula is as follows:
$W_{k/r,d/r,r,r}$ $=Reshape(W_{d,k})$ (4)
$bone_{d/r,r,r}$ $=Reshape(bone_{r,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/r,d/r,r,r}bone_{d/r,r,r}+bone_{d/r,r,r})$
It is evident that Bat can be defined in many other ways, as illustrated in the figure 4. In Section 6, we compared the performance of different partitioning strategies for the structure. The formula for Bone-row is as follows:
$W_{d/r,k/r,r,r}$ $=Reshape(W_{d,k})$ (5)
$bone_{k/r,r,r}$ $=Reshape(bone_{r,k})$
$\Delta W_{d,k}$ $=Reshape(W_{d/r,k/r,r,r}bone_{k/r,r,r}+bone_{k/r,r,r})$

### v7

# DiSHA: Dimension-Sharding Adaptation of Large Language Models with Fast Convergence and Fast Computation

@@ Abstract
Low-Rank Adaptation (LoRA), a prominent technique within the framework of Parameter-Efficient Fine-Tuning (PEFT), efficiently reduces the computational burden associated with adapting Large Language Models (LLMs) to downstream tasks, thereby enabling resource-constrained fine-tuning. However, existing researches have shown that LoRA suffers from slow convergence. To address this limitation, we introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to even fewer trainable parameters and faster convergence. Within DiSHA’s design space, we propose Block Affine Efficient Computation (Bone), a computationally efficient structure that delivers both high performance and efficiency. While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation (Bat), a nonlinear variant of DiSHA. Bat introduces nonlinearity by combining trainable matrices with original weight shards in a nonlinear manner, inducing nonlinearity in matrix updates without introducing additional parameters. Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both Natural Language Understanding and Natural Language Generation tasks, with significantly improved computational efficiency. Further analysis demonstrates that BAT enhances model capabilities by leveraging its nonlinear design.
[Uncaptioned image]https://github.com/huggingface/peft
Machine Learning, ICML
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  (Radford et al., 2019; Raffel et al., 2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is computationally costly. As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)(Xu et al., 2023) techniques and quantization methods have emerged to reduce the training costs of large models. Low-Rank Adaptation(LoRA) (Hu et al., 2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning. The figure  3 illustrates the structure visualization.
@@
However, extensive experiments (Ding et al., 2023; Liu et al., 2024b; Biderman et al., 2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ (Hayou et al., 2024), PISSA (Meng et al., 2024), and LoRA-GA(Wang et al., 2024). These methods have all brought unexpected improvements. By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
Figure 2: partition of weight matrix $\mathbf{W_{0}}$.
@@
Based on the aforementioned conjecture, we first propose a framework with a simpler update mechanism, called Dimension-Sharding Adaptation (DiSHA), designed to achieve larger initial gradients with fewer trainable parameters, thereby accelerating the convergence process. In DiSHA, the pre-trained weight matrix $\mathbf{W_{0}}$ is frozen and partitioned into multiple shard matrices during training, and all shard matrices share a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters. The design space of DiSHA is extensive, permitting the division of pre-trained weights matrix into dimensions of arbitrary magnitudes, as illustrated in Figure 2. Computational efficiencies vary among design choices, so we proposed an efficient and unified structure named Block Affine Efficient Computation (Bone) within a subset of the DiSHA design space and validated its effectiveness through extensive experiments. As the experiments progressed, we discovered that when fine-tuning with Bone, the updates between different shards within the same matrix are collinear; therefore, to introduce nonlinear updates, we proposed the Block Affine Transformation (Bat) method. Inspired by methods like PiSSA that leverage pre-trained weight matrix information, the updates between different shard matrices are no longer directly added to the trainable matrix. Instead, they need to be linearly transformed with the corresponding shard matrix before being summed. This approach not only leverages the information from the pre-trained weight matrices but also ensures that the updates between different shard matrices become nonlinear.
Our extensive evaluation shows that Bone demonstrates excellent performance on both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks, significantly surpassing LoRA and its variants. Moreover, as shown in Figure  5, Bone also achieves a significantly faster convergence rate compared to PiSSA. This demonstrates the feasibility of the DiSHA framework. Furthermore, Bone’s unique design not only resolves the unification challenges of the DiSHA framework, allowing seamless adaptation across different LLM architectures such as LLaMA, RWKV, and Mistral but also enables it to significantly outperform LoRA in both computational efficiency and memory usage. In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 (Xu et al., 2023), RWKV6 (Peng et al., 2024)). Based on the comparative experiments between Bone and Bat, it is evident that the non-linear updates of weight shards enhance the model’s performance ceiling.
@@
2. We proposed an efficient and adaptable structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
3. We propose Block Affine Transformation (Bat), which introduces shard-specific nonlinearity by integrating pre-trained weight information, without adding new parameters: A breakthrough to overcome the collinear update bottleneck inherent in both Bone.
@@ 2 Related Works
The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuning(Houlsby et al., 2019; He et al., 2022; Wang et al., 2022; Pfeiffer et al., 2020), prefix tuning(Liu et al., 2023; Li & Liang, 2021), prompt tuning(Brown, 2020; Liu et al., 2023; Lester et al., 2021; Razdaibiedina et al., 2023; Li & Liang, 2021), LoRA(Hu et al., 2021; Meng et al., 2024; Wang et al., 2024; Si et al., 2024), and layer-freezing methods such as LISA.
@@
LoRA, as one of the most popular PEFT techniques, has demonstrated that the updates to the weights exhibit a low intrinsic rank during adaptation. In LoRA, for a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $\mathbf{W}_{0}+\mathbf{\Delta W}=\mathbf{W}_{0}+\mathbf{BA}$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, $\mathbf{W}_{0}$ is frozen and does not receive gradient updates, while $\mathbf{A}$ and $\mathbf{B}$ contain trainable parameters. Note both $\mathbf{W}_{0}$ and $\mathbf{\Delta W}=\mathbf{BA}$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $h=\mathbf{W}_{0}x$, our modified forward pass yields:
$h=\mathbf{W}_{0}x+\mathbf{\Delta W}x=\mathbf{W}_{0}x+\mathbf{BA}x$ (1)
@@
Since A and B are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process. As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial learning rat, among other approaches.
2.1 Adjusting the initial learning rate of LoRA
LoRA+ extended this method by introducing independent learning rates for matrices $\mathbf{A}$ and $\mathbf{B}$ with a fixed ratio, improving the method’s efficiency. DoRA (Liu et al., 2024b) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately. This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning.
2.2 Improving the Initialization of LoRA
@@
PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA (Büyükakyüz, 2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\mathbf{A}$ and $\mathbf{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning.
Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
@@ 3 Method
3.1 Motivation: Issues with LoRA
We have identified two main limitations of LoRA and its variants: Firstly, LoRA with simple initialization suffers from slow convergence.
Secondly, while LoRA variants accelerate convergence, they also introduce many more complex operations.
Specifically, methods like PiSSA and OLoRA, which set the initialization by decomposing the pre-trained weights, increase the initialization time. Moreover, the initialization time grows as the model’s parameter size increases.
@@ 3.2 DiSHA: Dimension-Sharding Adaptation
To address these limitations and based on the conclusions from the gradient norm experiments A.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower trainable parameters and faster convergence by default. For a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $\mathbf{W}=\mathbf{W}_{0}+\mathbf{\Delta W}=\mathbf{W}_{0}+\operatorname{expand}(\mathbf{D})$, where $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}},\operatorname{expand}(\mathbf{D})\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$. During training, $\mathbf{W}_{0}$ is frozen and does not receive gradient updates, while the trainable matrix $\mathbf{D}$ is initialized to zero, ensuring $\operatorname{expand}(\mathbf{D})$ initially contributes no update to $\mathbf{W}_{0}$. Note both $\mathbf{W}_{0}$ and $\mathbf{\Delta W}=\operatorname{expand}(\mathbf{D})$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $y=\mathbf{W}_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
$y=\mathbf{W}_{0}x+\mathbf{\Delta W}x=\mathbf{W}_{0}x+\operatorname{expand}(\mathbf{D})x$ (2)
Here, we explain the concept of $\operatorname{expand}(\mathbf{D})$. As shown in the DiSHA structure in Figure  3, we partition $\mathbf{W}_{0}$ into multiple shards along the dimension of $\mathbf{D}$. The $\mathbf{D}$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $\mathbf{W}_{0}$. Finally, the updated shards are summed to update $\mathbf{W}_{0}$.
@@ 3.3 DiSHA Design Space
The design space of DiSHA is determined by its partitioned low-rank update mechanism. Given a pre-trained weight $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, the trainable matrix $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ governs updates through two configurable degrees of freedom:
1. Rank Selection: The ranks $r_{1}$ and $r_{2}$ independently control the parameterization complexity. For example:
• A wide matrix ($d\ll k$) benefits from $r_{1}\ll r_{2}$ to prioritize input-dimension compactness.
• A tall matrix ($d\gg k$) favors $r_{1}\gg r_{2}$ to minimize memory overhead.
2. Partitioning Strategy: The output dimension $d$ can be partitioned into $N$ shards with sizes $\{s_{1},...,s_{N}\}$, where $\sum_{i=1}^{N}s_{i}=d$. Each shard’s update is scaled by a row of $\mathbf{D}$:
$\operatorname{expand}(\mathbf{D})=\bigoplus_{i=1}^{N}\left(\mathbf{1}_{s_{i}}\otimes\mathbf{D}_{i}\right)\in\mathbb{R}^{d\times k}$
Here, $\mathbf{D}_{i}\in\mathbb{R}^{1\times r_{2}}$ is replicated $s_{i}$ times vertically.
As shown in Table 2, DiSHA can set fewer trainable parameters while supporting non-uniform partitioning (e.g., grouping attention heads in transformers). This allows for a wide design space in DiSHA.
3.4 Bone: Block Affine Efficient Computation
@@
In the previous section, we introduced the DiSHA framework, a method that significantly reduces the number of trainable parameters in the model. Due to the vast design space of the DiSHA framework, it is difficult for users to accurately determine suitable values for training. A good setting of $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks. For example, with the rise of technologies like GQA (Ainslie et al., 2023) and MQA (Shazeer, 2019), the matrix dimensions of self-attention modules ($\mathbf{W}_{k},\mathbf{W}_{v}$) vary greatly among different LLM models.
To enhance the adaptability and efficiency of DiSHA, we proposed Block Affine Efficient Computation (Bone), which reduces the memory footprint of intermediate parameters while enhancing computational efficiency through an innovative input aggregation mechanism.
Core Innovation: Blockwise Input Aggregation Bone redefines the trainable matrix as $\mathbf{D}\in\mathbb{R}^{r\times d}$ with a single rank parameter $r$, enabling unified adaptation across heterogeneous layers. The key insight is to decouple the input dimension $k$ from the parameterization by introducing blockwise summation:
1. Input Partitioning: For input $x\in\mathbb{R}^{b\times l\times k}$, partition the $k$-dimension into $r$ contiguous blocks of size $g=\lfloor k/r\rfloor$:
$x\rightarrow[x^{(1)},\dots,x^{(r)}],\quad x^{(i)}\in\mathbb{R}^{b\times l\times g}$
2. Blockwise Summation: Sum each partitioned block along the $k$-dimension:
$\mathbf{S}=\left[\sum_{j=1}^{g}x^{(1)}_{:,:,j},\dots,\sum_{j=1}^{g}x^{(r)}_{:,:,j}\right]\in\mathbb{R}^{b\times l\times r}$
3. Affine Projection: Compute the low-rank update via matrix multiplication with $\mathbf{D}$:
$\mathbf{\Delta W}x=\mathbf{D}^{\top}\mathbf{S}\quad(\mathbf{D}\in\mathbb{R}^{r\times d},\mathbf{S}\in\mathbb{R}^{b\times l\times r})$
The final forward pass becomes:
$y=\mathbf{W}_{0}x+\mathbf{D}^{\top}\mathbf{S}$
where $\mathbf{D}^{\top}\mathbf{S}\in\mathbb{R}^{b\times l\times d}$ matches the output dimension of $\mathbf{W}_{0}x$.
Mathematical Equivalence to DiSHA Bone’s blockwise summation is equivalent to DiSHA’s $\operatorname{expand}(\mathbf{D})$ operation under structured sparsity. Let $\operatorname{expand}(\mathbf{D})\in\mathbb{R}^{d\times k}$ be constructed by tiling each row of $\mathbf{D}$ over $g$ columns:
$\operatorname{expand}(\mathbf{D})_{i,j}=\mathbf{D}_{i,\lceil j/g\rceil}$
Then:
$\operatorname{expand}(\mathbf{D})x=\sum_{i=1}^{r}\mathbf{D}_{i}\sum_{j=1}^{g}x^{(i)}_{:,:,j}=\mathbf{D}^{\top}\mathbf{S}$
This equivalence allows Bone to avoid explicitly materializing $\operatorname{expand}(\mathbf{D})$, reducing memory from $O(dk)$ to $O(dr)$, and reducing flops from $O(bldk)$ to $O(blr(d+\frac{k}{r}))$